# Improving Gloss-free Sign Language Translation by Reducing Representation Density

### v1

@@ 1 Introduction
Sign languages are the primary form of communication for millions of deaf individuals. Sign language translation (SLT) aims to convert sign language into fluent spoken language sentences, which is a challenging task as it needs to extract information from continuous video and translate it into discrete text tokens. Most prior studies promoted the SLT by utilizing intermediate representations, namely gloss annotations, either directly or indirectly Camgoz et al., (2018); Yin and Read, (2020); Zhou et al., 2021a ; Chen et al., (2022); Zhang et al., 2023a ; Ye et al., 2023b . Gloss annotations are beneficial as they provide a simplified representation and sequential ordering of each gesture within continuous sign videos, which aids in representation learning for visual encoders. However, the creation of sign language translation datasets with gloss annotations is both resource-intensive and time-consuming.
@@
In this paper, we identify a representation density problem in sign language translation: the visual representations of sign gestures with distinct semantics are likely to be close in representation space. This problem is attributed to the nature of sign language, a form of visual language that utilizes intricate hand gestures, facial expressions, and body movements to convey the signer‘s message Pizzuto, (2003); Vinson et al., (2010). For example, in Figure 1, the signer performs sign gestures for opposite meanings, “RECIPROCATE” and “REVENGE”, with similar visual information (i.e., only subtle differences in facial movements). The visual encoder in SLT models will encode similar visual information to visual representations in adjacent representation space, even though they have distinct semantics. Without explicit gloss annotations, SLT models struggle to learn semantic boundaries in continuous sign videos and capture distinguishing visual representations for different sign gestures. As a result, the representation density problem poses a significant challenge for the SLT models in distinguishing between various sign gestures, leading to sharp performance drops. (Section  2.2).
@@
Further, we investigate various popular sign feature extraction methods, including gloss-based Camgoz et al., (2018); Min et al., (2021) and gloss-free Yin et al., (2023); Zhou et al., (2023), to systematically study the representation density problem. As shown in Figure 2, our investigation reveals that the representation density problem is prevalent across sign feature extraction methods. Specifically, due to the lack of gloss annotations, the representation density problem appears to be more serious in gloss-free methods. Then, we conduct extensive SLT experiments and observe that SLT models using gloss-free sign features as input consistently suffer a drop in performance in both sign language recognition and translation tasks compared to those using gloss-based sign features (Section  2.3). Therefore, we demonstrate that the representation density problem can be a bottleneck in restricting the improvement of gloss-free sign language translation.
More importantly, we propose a simple but effective contrastive learning strategy named SignCL to address the representation density problem. Specifically, SignCL draws the visual representations of sign gestures with identical semantics closer together and pushes those with different semantics farther apart. Experimental results show that SignCL can learn more distinctive feature representations and lead to significant improvements in terms of BLEU score on various well-known SLT frameworks (Section  4). To summarize, the main contributions of this work are as follows:
@@
  • We propose a simple but effective contrastive learning strategy, namely SignCL, to address the representation density problem. Our experiments demonstrate that SignCL can significantly enhance various well-known SLT frameworks. Specifically, SignCL yields a 39.01% BLEU score improvement for the Sign Language Transformer Camgoz et al., 2020a and a 45.58% increase for GFSLT-VLP Zhou et al., 2021a on the CSL-Daily dataset.
@@ 2 Representation Density Problem
This section investigates the representation density problem within existing sign feature extraction techniques and examines whether representation density influences sign language recognition and translation performance.
@@ 2.1 Preliminaries
  • Sign Recognition Pretrained (SRP) Camgoz et al., (2018): This approach leverages the sign language recognition datasets to train sign language recognition models and uses it as the feature extractor for the SLT task. Notably, the features released by Camgoz et al., (2018) have been widely adopted as input features in a range of works Camgoz et al., 2020b ; Zhou et al., 2021a ; Jin and Zhao, (2021); Yao et al., (2023); Ye et al., 2023b .
@@
$SDR(G_{i})=\frac{{D^{intra}_{G_{i}}}\,}{{avg.}D^{inter}_{G_{i}}}=\frac{D(G_{i})}{{Mean}_{\begin{subarray}{c}j\neq i\end{subarray}}\left(D(G_{i},G_{j})\right)}$ (1)
@@ 2.2 Demonstrating Representation Density Problem
Figure 3: Comparative analysis of representation density and its impact on sign language recognition (SLR) and translation (SLT). The left panel (a) shows the correlation between representation density and SLR accuracy across different sign feature types and sign gesture groups. The right panel (b) illustrates the performance drops in SLT caused by the representation density problem.
@@ 2.3 Demonstrating Performance Drop
(1) Sign Language Recognition: To evaluate the ability of the extracted sign features to distinguish between different semantic gestures, we use the NSLT Camgoz et al., 2020a to perform sign language recognition (SLR) tasks with various types of sign features Camgoz et al., (2018) as model input. Due to the limited number of samples for each gesture in the dev set, we rank the sign glosses based on their density using SDR$(G_{i})$ under SMKD features (see Eqn.  1). These glosses are then divided into nine groups (bins), each containing approximately 60 glosses. The average SDR$(G_{i})$ and recognition accuracy for each bin represents the overall density and mean accuracy of the glosses within that bin, respectively. (2) Sign Language Translation: This evaluation aims to demonstrate the impact of representation density on translation tasks. We evaluate various sign features as inputs to the Sign Language Transformer, including SRP, SMKD, I3D, VLP, and VLP+SignCL. These inputs are tested across different translation frameworks, such as NSLT Camgoz et al., (2018), Joint-SLT Camgoz et al., 2020a , and NSLT+SignCL. NSLT means use NSLT to perform SLT without CTC loss (gloss-free) and the NSLT+SignCL configuration integrates the proposed contrastive learning strategy into the encoder of NSLT Camgoz et al., (2018) models, as detailed in Section 3.
@@
  • Performance suffers from representation density. We consistently observed a negative relationship between representation density and performance across all feature types and tasks. Higher representation density leads to worse accuracy in SLR and lower BLEU scores in SLT. Specifically, an increase in the representation density ratio by 26% can result in a 36% performance drop in NSLT.
@@ 3.1 Sign Contrastive Learning
$\begin{cases}\text{positive pair }(f_{st},f_{ed}^{+})\text{:}&|f_{ed}^{+}-f_{st}|\leq 1\\ \text{negative pair }(f_{st},f_{ed}^{-})\text{:}&|f_{ed}^{-}-f_{st}|>2*margin\\ \end{cases}$ (4)
$\mathcal{L}_{SignCL}=\frac{1}{N}\sum_{{st}=1}^{N}\left[\text{d}(f_{st},f_{ed}^{+})+\max(0,m-\text{d}(f_{st},f_{ed}^{-}))\right]$ (5)
@@ 3.2 Integrating Contrastive Learning into Sign Language Translation Tranining
$\mathcal{L}=\lambda*{\mathcal{L}}_{\text{SignCL}}+{\mathcal{L}}_{MLE}$ (6)
@@ 4.1 Experiments on Sign Language Transformer
In Section 2, we investigate the representation density problem using the Sign Language Transformer (NSLT)Camgoz et al., 2020a and the PHOENIX-2014T DatasetCamgoz et al., (2018). These benchmarks are chosen for their established relevance in sign language translation research, including gloss-based and gloss-free based. Here, we first conduct experiments on the same framework and dataset to facilitate direct comparison with the prior empirical analyses.
Experiment Settings: In this experiment, we introduce SignCL as additional supervision information in the encoder of NSLT under gloss-free settings. This enhanced model is referred to as NSLT+SignCL.
Results and Findings: As shown in Figure 3(a) and Figure 3(b), integrating SignCL into the NSLT has demonstrated significant improvements in reducing representation density and improving translation performance across all test conditions. Notably, the visualization results demonstrate that SignCL encourages NSLT to learn a more distinguishing feature distribution and reduces the Sign Density Ratio (SDR) of the baseline NSLT from 92.59% to 81.30%.
Figures 3(a) and 3(b) show the experiments on SLR and SLT tasks using features with different SDRs as inputs to NSLT. The results show that the representation density reduction has led to observable improvements in both recognition accuracy (red line vs. purple line in Figure 3(a)) and translation BLEU score (purple point vs. red point in Figure 3(b)). More specific numbers and experiments on the CSL-Daily dataset are listed in the Appendix A.3.3.
@@ 4.2 Experiments on Gloss-free Sign Language Translation
Results and Findings: Tables 1 and 2 compare our proposed methods with existing gloss-free sign language translation approaches. The results demonstrate that integrating the proposed SignCL strategy into the GFSLT-VLP framework consistently reduces representation density and significantly boosts translation performance, whether SignCL is applied during pretraining, finetuning, or both stages. Specifically, compared to the baseline model GFSLT-VLP  Zhou et al., (2023), our approach achieves a substantial improvement of 45.58% (+5.06) in the BLEU-4 score on the CSL-Daily dataset, without any increase in the number of parameters. Additionally, despite having significantly fewer parameters ($\sim$600M vs. $\sim$1.7B), our approach achieves better performance than Sign2GPT Wong et al., (2024), which leverages large-scale pretrained vision and language models for sign language translation.
@@ 4.3 Qualitative Analysis
5 Related Works
5.1 Sign Language Translation
Sign Language Translation (SLT) methods can be broadly categorized into gloss-based and gloss-free approaches. For gloss-based methods, an essential factor is to directly or indirectly employ sign gloss annotations to improve sign video encoder performance Camgoz et al., (2018); Yin and Read, (2020); Zhou et al., 2021a ; Chen et al., (2022); Zhang et al., 2023a . These methods often employ Connectionist Temporal Classification Graves et al., (2006) (CTC) loss to perform sign language recognition Camgoz et al., 2020a . Joint-SLT Camgoz et al., 2020a firstly introduces a multitask encoder-decoder framework with a CTCloss to soft-match sign representations and gloss sequences. STMC-T Zhou et al., 2021b introducing intra-cue and inter-cue CTC loss to model multi-cue sequence information. Despite their effectiveness, creating SLT datasets with gloss annotations is resource-intensive and time-consuming. Gloss-free methods have emerged as a promising alternative, as they do not rely on gloss annotations during training, making them more generalizable. And recently, a growing body of literature has promoted the gloss-free SLT, such as GASLT Yin et al., (2023) proposed local gloss attention to mimic gloss assistant, GFSLT Zhou et al., (2023) adapted CLIP to do visual-language pretraining, and Sign2GPT Wong et al., (2024) promoted performance by making use of large-scale pre-trained vision and language models. Nonetheless, the performance of gloss-free methods still significantly lags behind that of gloss-based approaches.
5.2 Contrastive Learning
Contrastive Learning Jaiswal et al., (2020), a popular unsupervised learning algorithm, aims to learn effective representations by pulling positive pairs closer together and pushing negative pairs farther apart. This approach has been widely utilized in both Natural Language Processing (NLP) and Computer Vision (CV). In Sign Language Translation (SLT), Jin and Zhao, (2021) utilize Contrastive Learning to create a Signer-Independent SLT model, using videos demonstrating signs from different signers as positive samples. Additionally, Gan et al., (2023) propose a visual-level contrastive learning method with various image augmentation strategies. ConSLT Fu et al., (2023) do contrastive learning for effective token representation learning in text decoder. Zhou et al., (2023) and Cheng et al., (2023) employ contrastive learning techniques to align video and text representations in SLT. In this paper, we are the first one to address the representation density problem, focusing particularly on visual gesture duration as a central aspect.
5.3 Representation Density
Representation Density is often a focal point in classification tasks, also known as category density Bengio et al., (2013); Xie et al., (2016); Schroff et al., (2015); Liu et al., (2016). This concept pertains to the compactness and clarity of feature representations across different categories. In the context of sign language, various methods have been developed to address the subtle nuances of sign actions. TSPNet Li et al., 2020b proposes a temporal hierarchical attention network to learn segmented representations. HST-GNN Kan et al., (2022) utilizes a hierarchical spatio-temporal graph neural network to learn graph representations from multiple perspectives. GLE-Net Hu et al., (2021) employs global contextual relationships and fine-grained cues to distinguish non-manual-aware features in isolated Sign Language Recognition. These methods are beneficial for addressing the subtleties of sign language movements. However, integrating them into existing state-of-the-art frameworks presents significant challenges, often resulting in performance disparities when compared to the SOTA. This paper is the first to propose the concept of representation density within this field and introduces SignCL, which enhances the current mainstream transformer-based frameworks.

### v2

@@ 1 Introduction
Sign languages are the primary form of communication for millions of deaf individuals. Sign language translation (SLT) aims to convert sign language into fluent spoken language sentences, which is a challenging task as it needs to extract information from continuous video and translate it into discrete text tokens. Most prior studies promoted the SLT by utilizing intermediate representations, namely gloss annotations, either directly or indirectly Camgoz et al., (2018); Yin and Read, (2020); Zhou et al., 2021a ; Chen et al., (2022); Zhang et al., 2023a ; Ye et al., 2023b ; Wang et al., (2024). Gloss annotations are beneficial as they provide a simplified representation and sequential ordering of each gesture within continuous sign videos, which aids in representation learning for visual encoders. However, the creation of sign language translation datasets with gloss annotations is both resource-intensive and time-consuming.
@@
In this paper, we identify a representation density problem in sign language translation: the visual representations of sign gestures with distinct semantics are likely to be close in representation space. This problem is attributed to the nature of sign language, a form of visual language that utilizes intricate hand gestures, facial expressions, and body movements to convey the signer‘s message Pizzuto, (2003); Vinson et al., (2010); Guo et al., (2022). For example, in Figure 1, the signer performs sign gestures for opposite meanings, “RECIPROCATE” and “REVENGE”, with similar visual information (i.e., only subtle differences in facial movements). The visual encoder in SLT models will encode similar visual information to visual representations in adjacent representation space, even though they have distinct semantics. Without explicit gloss annotations, SLT models struggle to learn semantic boundaries in continuous sign videos and capture distinguishing visual representations for different sign gestures. As a result, the representation density problem poses a significant challenge for the SLT models in distinguishing between various sign gestures, leading to sharp performance drops. (Section  3.2).
@@
Further, we investigate various popular sign feature extraction methods, including gloss-based Camgoz et al., (2018); Min et al., (2021) and gloss-free Yin et al., (2023); Zhou et al., (2023), to systematically study the representation density problem. As shown in Figure 2, our investigation reveals that the representation density problem is prevalent across sign feature extraction methods. Specifically, due to the lack of gloss annotations, the representation density problem appears to be more serious in gloss-free methods. Then, we conduct extensive SLT experiments and observe that SLT models using gloss-free sign features as input consistently suffer a drop in performance in both sign language recognition and translation tasks compared to those using gloss-based sign features (Section  3.3). Therefore, we demonstrate that the representation density problem can be a bottleneck in restricting the improvement of gloss-free sign language translation.
More importantly, we propose a simple but effective contrastive learning strategy named SignCL to address the representation density problem. Specifically, SignCL draws the visual representations of sign gestures with identical semantics closer together and pushes those with different semantics farther apart. Experimental results show that SignCL can learn more distinctive feature representations and lead to significant improvements in terms of BLEU score on various well-known SLT frameworks (Section  5). To summarize, the main contributions of this work are as follows:
@@
  • We propose a simple but effective contrastive learning strategy, namely SignCL, to address the representation density problem. Our experiments demonstrate that SignCL can significantly enhance various well-known SLT frameworks. Specifically, SignCL yields a 39% BLEU score improvement for the Sign Language Transformer Camgoz et al., 2020a and a 46% BLEU increase for GFSLT-VLP Zhou et al., 2021a on the CSL-Daily dataset.
2 Related Works
2.1 Sign Language Translation
Sign Language Translation (SLT) methods can be broadly categorized into gloss-based and gloss-free approaches. For gloss-based methods, an essential factor is to directly or indirectly employ sign gloss annotations to improve sign video encoder performance Camgoz et al., (2018); Yin and Read, (2020); Zhou et al., 2021a ; Chen et al., (2022); Zhang et al., 2023a ; Chen et al., 2021a . These methods often employ Connectionist Temporal Classification Graves et al., (2006) (CTC) loss to perform sign language recognition Camgoz et al., 2020a . Joint-SLT Camgoz et al., 2020a firstly introduces a multitask encoder-decoder framework with a CTCloss to soft-match sign representations and gloss sequences. STMC-T Zhou et al., 2021b introducing intra-cue and inter-cue CTC loss to model multi-cue sequence information. Despite their effectiveness, creating SLT datasets with gloss annotations is resource-intensive and time-consuming. Gloss-free methods have emerged as a promising alternative, as they do not rely on gloss annotations during training, making them more generalizable. And recently, a growing body of literature has promoted the gloss-free SLT, such as GASLT Yin et al., (2023) proposed local gloss attention to mimic gloss assistant, GFSLT Zhou et al., (2023) adapted CLIP to do visual-language pretraining, and Sign2GPT Wong et al., (2024) promoted performance by making use of large-scale pre-trained vision and language models. Nonetheless, the performance of gloss-free methods still significantly lags behind that of gloss-based approaches.
2.2 Contrastive Learning
Contrastive Learning Jaiswal et al., (2020); Zhou et al., (2024); Lin et al., (2024), a popular unsupervised learning algorithm, aims to learn effective representations by pulling positive pairs closer together and pushing negative pairs farther apart. This approach has been widely utilized in both Natural Language Processing and Computer Vision Dai et al., (2023). In Sign Language Translation (SLT), Jin and Zhao, (2021) utilize Contrastive Learning to create a Signer-Independent SLT model, using videos demonstrating signs from different signers as positive samples. Additionally, Gan et al., (2023) proposes a visual-level contrastive learning method with various image augmentation strategies. ConSLT Fu et al., (2023) do contrastive learning for effective token representation learning in text decoder. Zhou et al., (2023) and Cheng et al., (2023) employ contrastive learning techniques to align video and text representations in SLT. In this paper, we are the first one to address the representation density problem, focusing particularly on visual gesture duration as a central aspect.
2.3 Representation Density
Representation Density is often a focal point in classification tasks, also known as category density Bengio et al., (2013); Xie et al., (2016); Schroff et al., (2015); Liu et al., (2016); Ma et al., (2021); Zhang et al., 2023c ; Lin et al., (2022); Dai et al., (2024). This concept pertains to the compactness and clarity of feature representations across different categories. In the context of sign language, various methods have been developed to address the subtle nuances of sign actions. TSPNet Li et al., 2020b proposes a temporal hierarchical attention network to learn segmented representations. HST-GNN Kan et al., (2022) utilizes a hierarchical spatio-temporal graph neural network to learn graph representations from multiple perspectives. GLE-Net Hu et al., (2021) employs global contextual relationships and fine-grained cues to distinguish non-manual-aware features in isolated Sign Language Recognition. These methods are beneficial for addressing the subtleties of sign language movements. However, integrating them into existing state-of-the-art frameworks presents significant challenges, often resulting in performance disparities when compared to the SOTA. This paper is the first to propose the concept of representation density within this field and introduces SignCL, which enhances the current mainstream transformer-based frameworks.
@@ 3 Representation Density Problem
This section investigates and identifies the representation density problem within existing sign feature extraction techniques, and examines whether representation density bottlenecks sign language recognition and translation performance.
@@ 3.1 Preliminaries
  • Sign Recognition Pretrained (SRP) Camgoz et al., (2018): This approach leverages the sign language recognition datasets to train sign language recognition models and uses it as the feature extractor for the SLT task. Notably, the features released by Camgoz et al., (2018) have been widely adopted as input features in a range of works Camgoz et al., 2020b ; Zhou et al., 2021a ; Jin and Zhao, (2021); Yao et al., (2023); Ye et al., 2023b ; Chen et al., 2021b .
@@
$SDR(G_{i})=\frac{{D^{intra}_{G_{i}}}\,}{{avg.}D^{inter}_{G_{i}}}=\frac{D(G_{i})}{{Mean}_{\begin{subarray}{c}j\neq i\end{subarray}}\left(D(G_{i},G_{j})\right)}.$ (1)
@@ 3.2 Demonstrating Representation Density Problem
(a,b,c) Gloss-free Feature Extraction           (d,e) Gloss-based Feature Extraction
@@
Figure 3: Comparative analysis of representation density and its impact on sign language recognition (SLR) and translation (SLT). The left panel (a) shows the correlation between representation density and SLR accuracy across different sign feature types and sign gesture groups. Binning in this context is based on sorting by gloss density within a group, where higher bins indicate higher density. The right panel (b) illustrates the performance drops in SLT caused by the representation density problem. This figure assesses both the recognition and translation accuracies, reflecting how denser representations impact these metrics.
@@ 3.3 Demonstrating Performance Drop
We set up tasks to examine whether representation density bottlenecks sign language recognition and translation performance.
  • Sign Language Recognition: To evaluate the ability of the extracted sign features to distinguish between different semantic gestures, we use the NSLT Camgoz et al., 2020a to perform sign language recognition (SLR) tasks with various types of sign features Camgoz et al., (2018) as model input. Due to the limited number of samples for each gesture in the dev set, we rank the sign glosses based on their density using SDR$(G_{i})$ under SMKD features (see Eqn.  1). These glosses are then divided into nine groups (bins), each containing approximately 60 glosses. The average SDR$(G_{i})$ and recognition accuracy for each bin represents the overall density and mean accuracy of the glosses within that bin, respectively.
  • Sign Language Translation: This evaluation aims to demonstrate the impact of representation density on translation tasks. We evaluate various sign features as inputs to the Sign Language Transformer, including SRP, SMKD, I3D, VLP, and VLP+SignCL. These inputs are tested across different translation frameworks, such as NSLT Camgoz et al., (2018), Joint-SLT Camgoz et al., 2020a , and NSLT+SignCL. NSLT means use NSLT to perform SLT without CTC loss (gloss-free) and the NSLT+SignCL configuration integrates the proposed contrastive learning strategy into the encoder of NSLT Camgoz et al., (2018) models, as detailed in Section 4.
@@
  • Performance suffers from representation density. We consistently observed a negative relationship between representation density and performance across all feature types and tasks. Higher representation density leads to worse accuracy in SLR and lower BLEU scores in SLT. Specifically, an increase in the representation density ratio by 26% can result in a 39% performance drop in NSLT.
@@ 4.1 Sign Contrastive Learning
$\begin{cases}\text{positive pair }(f_{st},f_{ed}^{+})\text{:}&|f_{ed}^{+}-f_{st}|\leq 1\\ \text{negative pair }(f_{st},f_{ed}^{-})\text{:}&|f_{ed}^{-}-f_{st}|>2*margin\\ \end{cases};$ (4)
$\mathcal{L}_{SignCL}=\frac{1}{N}\sum_{{st}=1}^{N}\left[\text{d}(f_{st},f_{ed}^{+})+\max(0,m-\text{d}(f_{st},f_{ed}^{-}))\right];$ (5)
@@ 4.2 Integrating Contrastive Learning into Sign Language Translation Tranining
$\mathcal{L}=\lambda*{\mathcal{L}}_{\text{SignCL}}+{\mathcal{L}}_{MLE};$ (6)
@@ 5.1 Experiments on Sign Language Transformer
In Section 3, we investigate the representation density problem using the Sign Language Transformer (SLT) Camgoz et al., 2020a and the PHOENIX-2014T Camgoz et al., (2018) and CSL-Daliy Zhou et al., 2021a Dataset. These benchmarks are chosen for their established relevance in sign language translation research, including gloss-based and gloss-free based. Here, we first conduct experiments on the same framework and dataset to facilitate direct comparison with the prior empirical analyses.
Experiment Settings: In this experiment, we introduce SignCL as additional supervision information in the encoder of SLT under gloss-free settings. This enhanced model is referred to as +SignCL.
Results and Findings: The integration of SignCL into the SLT has significantly improved translation performance across all test conditions by reducing the representation density, as shown in Table 1. Notably, SignCL encourages SLT to learn a more distinct feature distribution, reducing the Sign Density Ratio (SDR) significantly, e.g., 66.23 to 62.18 and 92.59 to 81.30.
Figures 3(a) and 3(b) show experiments on SLR and SLT tasks using features with varying SDRs as inputs to SLT. The representation density reduction leads to observable improvements in both recognition accuracy (red line vs. purple line in Figure 3(a)) and translation BLEU score (purple point vs. red point in Figure 3(b)). Further details and additional experiment results on the CSL-Daily dataset are provided in Appendix A.4.3.
Table 1 presents a comparative analysis of representation density and performance on the PHOENIX-2014T dataset. The inclusion of SignCL during VLP feature extraction or SLT training processes significantly enhances performance metrics. WERs (Word Error Rates) in the gloss-free set, derived from an independent SLR task, are specifically used to probe the quality of sign features and do not participate in the SLT training process. This analysis underscores the significant enhancements brought by SignCL in terms of both efficiency and effectiveness in SLT frameworks.
Feature Type PHOENIX-2014T CSL-Daily
[…]
                                                  Gloss-based
Joint-SLT Camgoz et al., 2020b / Self-Mutual KD Min et al., (2021) 66.23 25.38 22.79 48.34 29.52 11.61
+ SignCL into Feature Extraction 62.18 24.76 23.23 - - -
+ SignCL into Finetuning 66.23 25.12 22.92 - - -
+ SignCL into both 62.18 24.58 23.46 - - -
                                                  Gloss-free
SLT Camgoz et al., (2018) / VLP Pretrained Zhou et al., (2023) 92.59 69.72 10.73 76.55 85.78 1.82
+ SignCL into Feature Extraction 81.30 63.33 12.04 68.39 80.71 2.15
+ SignCL into Finetuning 92.59 - 12.14 76.55 - 2.19
+ SignCL into both 81.30 - 13.51 68.39 - 2.53
Table 1: Comparative analysis of representation density and performance on the PHOENIX-2014T dataset. "+SignCL" indicates the inclusion of the proposed contrastive learning strategy during VLP (Video Language Processing) feature extraction or SLT (Sign Language Translation) training processes. WERs (Word Error Rates) in the gloss-free set are derived from an independent SLR (Sign Language Recognition) task, used specifically for probing the quality of sign features. These WERs do not participate in the SLT training process.
@@ 5.2 Experiments on Gloss-free Sign Language Translation
Results and Findings: Tables 2 and 3 compare our proposed methods with existing gloss-free sign language translation approaches. The results demonstrate that integrating the proposed SignCL strategy into the GFSLT-VLP framework consistently reduces representation density and significantly boosts translation performance, whether SignCL is applied during pretraining, finetuning, or both stages. Specifically, compared to the baseline model GFSLT-VLP  Zhou et al., (2023), our approach achieves a substantial improvement of 45.58% (+5.06) in the BLEU-4 score on the CSL-Daily dataset, without any increase in the number of parameters. Additionally, despite having significantly fewer parameters ($\sim$600M vs. $\sim$1.7B), our approach achieves better performance than Sign2GPT Wong et al., (2024), which leverages large-scale pretrained vision and language models for sign language translation.